Two identical small balls A and B each of mass m connected by a light inextensible string of length l are placed on a frictionless horizontal floor. With what velocity u must the ball B be projected vertically upwards so that the ball A leaves the floor?

Grade:12

3 Answers

Sanyam Shah
3 years ago
Apply the equation when system becomes vertical:
T=mg (to lift it from the ground)
T=mu2/l (centripetal force provided by tension to the other ball)
Solve them and get answer u=root(gl)
Sanyam Shah
3 years ago
Above solution is wrong
u = root(gl) is the final velocity of both the balls performing circular motion at that point to find initial speed u
apply energy cnservation law:
1/2mu02 = 1/2mu2 + mgl 
now keep the value of u in above equation and get answer u>=root(3gl)
 
Karl Marx
2 years ago
When Ball A leaves floor, N=0
So, T = mg      ....(i)
Also since Ball B undergoes circular motion wrt Ball A 
So, T = mv^2/l   ....(ii)
 Now conserving Mechanical energy at ground and highest point, 
 1/2mu^2 = 1/2 mv^2 + mgl    ….(iii)
Solving all three equations we get,
u = root(3gl) [And thats the final correct answer].